# Missing Metals in DQ Stars; a Compelling Clue to their Origin††thanks: Dedicated to the memory of Prof. G. Fontaine (1948–2019)

### v2

@@ 4.5 Contrasts with DZ and DC stars
The most obvious features of the cumulative mass histogram is that DQ white dwarfs have a larger fraction of lower masses, the DC stars have a notable fraction of higher masses, while the planetary system hosting DZ stars are somewhat in the middle of these two. Another feature is the concentration of DQ white dwarf masses in the range M, shown by a steepness in the distribution that is not observed for DZ or DC stars.
@@
The tangential velocity distributions tell a similar story but with less distinction between the DZ and DC white dwarfs. The DZ stars appear to include few stars with relatively high velocities, consistent with planetary hosts coming from a primarily young disk population, where star formation likely benefitted from a comparatively metal-rich environment. The DQ stars are modest outliers compared to their helium atmosphere, DZ and DC white dwarf counterparts, and seem to be missing stars at the lowest velocities, below around 20 km s. This may imply the DQ white dwarfs are somewhat older than the DZ and DC stars.
Another possibility is that DQ white dwarfs have received significant velocity kicks from asymmetric mass loss, and which was not experienced by the DZ and DC stars. Only modest white dwarf recoil, no larger than 1 km s, has been inferred using a large and pristine catalog of wide Gaia binaries, via comparison of those pairs with two main-sequence stars, versus those containing at least one white dwarf (El-Badry & Rix, 2018). However, based on the large separation between binary components in that study, this modest recoil may apply only to stars that evolved in effective isolation. In contrast, close binary evolution may induce larger kicks that may account for the lack of low space velocities in the DQ stars.
@@ 5.5 Progenitors of DQ stars have high-intermediate masses
A plausible theory for the lack of planetary material in and around DQ white dwarfs is inhibited planet formation, and one possibility is via higher-mass main-sequence stars. If DQ white dwarfs are the progeny of stars with masses greater than M, then it could be argued that this is consistent with their lack of circumstellar matter, and pollution from planetary debris. Around such luminous stars, protoplanetary disk lifetimes may be insufficient to form planetary precursors and full-fledged planets (Kennedy & Kenyon, 2009; Kunitomo et al., 2021), and there are currently few confirmed cases of planets detected towards stars with masses above 3 M (e.g. Reffert et al. 2015). However, the expectations of this hypothesis would be that the remnant masses of DQ white dwarfs would be higher than the field, not lower as observed and modeled. And while such progeny may be prone to detectable magnetism (Caiazzo et al., 2020), they should be a relatively young kinematical population, having spent little time on the main sequence. These predictions fail to match the properties of DQ stars as shown in Figure 6. Furthermore, higher-mass parent stars would not address the lack of post-common envelope binaries, unless it were shown that this leads more commonly to companion destruction during the post-main sequence.
@@ 6 A DQ binary population outlook
For context, the solar zodiacal cloud is replenished at a rate in the range g s (Nesvorný et al., 2011; Rigley & Wyatt, 2022), and the rare DQZ exhibit accretion rates lower, not to mention the typical DQ upper limits at least lower, which approaches the accretion rate of extraterrestrial material onto the Earth (Love & Brownlee, 1993; Yada et al., 2004). It may be possible to detect multiple heavy elements in the rare DQZ stars, and if successful, their ratios can be compared to the planetary abundances associated with DAZ, DBZ, and DZ white dwarfs. While challenging owing to the low abundances and line strengths, it may provide an important clue to the nature of these stars, which are now, hopefully, less of an astrophysical enigma.
@@ Acknowledgements
J. Farihi thanks D. Koester and G. Fontaine for key discussions over the years that helped to form the foundation of this work, E. Dennihy for confirming the fact that no DQ star has a suspected infrared excess from circumstellar dust, A. Rebassa-Mansergas for sharing his full white dwarf-main sequence binary catalog, and valuable exchanges with E. B. Bauer, R. R. Rafikov, and J. J. Eldridge. Several colleagues provided feedback on an earlier version of the manuscript, including those mentioned above, as well as T. von Hippel, A. J. Mustill, A. Swan, and B. Zuckerman. The authors acknowledge the European Southern Observatory for the award of telescope time via programs 095.D-0706, 096.D-0076, and 097.D-0063. This work has made use of the ESA Gaia mission, processed by the Gaia DPAC. J. Farihi acknowledges support from STFC grant ST/R000476/1. T. G. Wilson acknowledges support from STFC consolidated grants ST/R000824/1, ST/V000861/1, and UKSA grant ST/R003203/1.

### v3

@@ 4.5 Contrasts with DZ and DC stars
The most obvious features of the cumulative mass histogram is that DQ white dwarfs have a larger fraction of lower masses, the DC stars have a notable fraction of higher masses, while the planetary system hosting DZ stars are somewhat in the middle of these two. Another feature is the concentration of DQ white dwarf masses in the range M⊙, shown by a steepness in the distribution that is not observed for DZ or DC stars.
@@
The tangential velocity distributions tell a similar story but with less distinction between the DZ and DC white dwarfs. The DZ stars appear to include few stars with relatively high velocities, consistent with planetary hosts coming from a primarily young disk population, where star formation likely benefitted from a comparatively metal-rich environment. The DQ stars are modest outliers compared to their helium atmosphere, DZ and DC white dwarf counterparts, and seem to be missing stars at the lowest velocities, below around 20 km s-1. This may imply the DQ white dwarfs are somewhat older than the DZ and DC stars.
Another possibility is that DQ white dwarfs have received significant velocity kicks from asymmetric mass loss, and which was not experienced by the DZ and DC stars. Only modest white dwarf recoil, no larger than 1 km s-1, has been inferred using a large and pristine catalog of wide Gaia binaries, via comparison of those pairs with two main-sequence stars, versus those containing at least one white dwarf (El-Badry & Rix, 2018). However, based on the large separation between binary components in that study, this modest recoil may apply only to stars that evolved in effective isolation. In contrast, close binary evolution may induce larger kicks that may account for the lack of low space velocities in the DQ stars.
@@ 5.5 Progenitors of DQ stars have high-intermediate masses
A plausible theory for the lack of planetary material in and around DQ white dwarfs is inhibited planet formation, and one possibility is via higher-mass main-sequence stars. If DQ white dwarfs are the progeny of stars with masses greater than M⊙, then it could be argued that this is consistent with their lack of circumstellar matter, and pollution from planetary debris. Around such luminous stars, protoplanetary disk lifetimes may be insufficient to form planetary precursors and full-fledged planets (Kennedy & Kenyon, 2009; Kunitomo et al., 2021), and there are currently few confirmed cases of planets detected towards stars with masses above 3 M⊙ (e.g. Reffert et al. 2015). However, the expectations of this hypothesis would be that the remnant masses of DQ white dwarfs would be higher than the field, not lower as observed and modeled. And while such progeny may be prone to detectable magnetism (Caiazzo et al., 2020), they should be a relatively young kinematical population, having spent little time on the main sequence. These predictions fail to match the properties of DQ stars as shown in Figure 6. Furthermore, higher-mass parent stars would not address the lack of post-common envelope binaries, unless it were shown that this leads more commonly to companion destruction during the post-main sequence.
@@ 6 A DQ binary population outlook
For context, the solar zodiacal cloud is replenished at a rate in the range g s-1 (Nesvorný et al., 2011; Rigley & Wyatt, 2022), and the rare DQZ exhibit accretion rates lower, not to mention the typical DQ upper limits at least lower, which approaches the accretion rate of extraterrestrial material onto the Earth (Love & Brownlee, 1993; Yada et al., 2004). It may be possible to detect multiple heavy elements in the rare DQZ stars, and if successful, their ratios can be compared to the planetary abundances associated with DAZ, DBZ, and DZ white dwarfs. While challenging owing to the low abundances and line strengths, it may provide an important clue to the nature of these stars, which are now, hopefully, less of an astrophysical enigma.
@@ Acknowledgements
The authors thank an anonymous reviewer for a careful reading of the manuscript. J. Farihi thanks D. Koester and G. Fontaine for key discussions over the years that helped to form the foundation of this work, E. Dennihy for confirming the fact that no DQ star has a suspected infrared excess from circumstellar dust, A. Rebassa-Mansergas for sharing his full white dwarf-main sequence binary catalog, and valuable exchanges with E. B. Bauer, R. R. Rafikov, and J. J. Eldridge. Several colleagues provided feedback on an earlier version of the manuscript, including those mentioned above, as well as T. von Hippel, A. J. Mustill, A. Swan, and B. Zuckerman. The authors acknowledge the European Southern Observatory for the award of telescope time via programs 095.D-0706, 096.D-0076, and 097.D-0063. This work has made use of the ESA Gaia mission, processed by the Gaia DPAC. J. Farihi acknowledges support from STFC grant ST/R000476/1. T. G. Wilson acknowledges support from STFC consolidated grants ST/R000824/1, ST/V000861/1, and UKSA grant ST/R003203/1.